A merchant purchased a line of handbags for R172,00 each. She wants to offer a 14% discount off the listed

selling price and still make a 24% profit. The price that she should mark on the label is


1
Expert's answer
2021-04-28T04:34:30-0400

"Cost \ of \ the \ bags \ is \ C=R172,000\\
She \ gained \ a \ profit \ of \ 24\%\Rightarrow Profit =172,000*\frac{24}{100}=1720*24=R" 41,280"Sale \ price = Cost +Profit\\
=R172,000+R41,280\\
=R213,280\\
Sale \ price = R 213,280\\
Discount \ given \ on \ sale \ price \ is 14\%\\
Let \ 'x' \ be \ the \ tag \ price .\\
Since \ the \ discount \ 14\% \ is \ given \ on \ x,\\
the \ sale \ price =\frac{(100-14)}{100}*x\\
\Rightarrow \frac{86}{100}*x=213,280\\
\Rightarrow x=\frac{(213280*100)}{86}=248,000\\
\therefore The \ price\ tag \ x=R 248,000"